Circular No. 8133

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 J1 (NEAT) K. J. Lawrence, Jet Propulsion Laboratory, reports the discovery by

NEAT of a comet (discovery observation below). Following posting on the NEO Confirmation Page, other CCD observers reported the following total magnitudes and coma diameters: May 14.5 UT, m_{1} = 16.4–17.0, 10′′ (P. Holvorcem, Tenagra II 0.81-m telescope; three co-added 120-s exposures); 15.5, 17.5, 8′′ (J. Young, Table Mountain, CA, 0.6-m reflector).

2003 UT  α_{2000}  δ_{2000}  m_{1}
May 13.59182  21^{h}00^{m}45.^{s}37  −15^{o}10′51′′.3  19.4

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-K15.

T = 2003 Sept.11.964 TT  ω = 192.^{o}477
Ω = 122.095 }2000.0
q = 5.19493 AU  i = 98.041

GRB 030329 AND SUPERNOVA 2003dh K. S. Kawabata, G. Kosugi, M. Iye, T. Sasaki, and Y. Ohyama, Na-

tional Astronomical Observatory of Japan; J. Deng and K. Nomoto, University of Tokyo; P. Mazzali, Trieste Observatory, report, on behalf of the Subaru Supernova/GRB ToO group, on optical spectropolarimetry (range 590–1020 nm) and spectroscopy (ranges 420–600 and 590–1020 nm) of the afterglow of GRB 030329 (IAUC 8101) and its supernova component SN 2003dh (IAUC 8108, 8114) obtained on May 8.3 and 9.3 UT with the Subaru 8.2-m telescope (+ FOCAS): “The spectrum was compared to those of other type-Ic hypernovae. The red spectrum resembles closely those of the hypernova SN 1997ef at 34 or 42 days after explosion (Mazzali et al. 2000, Ap.J. 545, 407). The match with SN 1998bw spectra is not as good between rest wavelengths 700 and 800 nm, but the spectra obtained ∼ 33–36 days after the outburst of GRB 980425 are preferred (Patat et al. 2001, Ap.J. 555, 900). The epochs are comparable to the rest-frame age of SN 2003dh at the time of the observations (days 34–35, assuming z = 0.1685; GCN 2020; IAUC 8108). The spectrum shows a significant excess in the blue, which may be a contribution of the γ-ray-burst afterglow and/or some other sources such as the host galaxy. Also, our preliminary results imply that polarization is not significant (< 1 percent), at least for rest wavelengths 500–600 nm. Detailed analysis is in progress.”

2003 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green